XYZ Corp. has $34 million dollars that it wishes to pay out as dividends, and a cost of equity of 0.22. If it pays out $12 million in dividends today, what dividend should they pay 1 year from today to make shareholders indifferent between this dividend scheme and one where shareholders get all of the cash today?
Answer #1

Dividend to be paid one year from today=(Cash today-Dividend today)*(1+cost of equity)=(34-12)*(1+0.22)=26.84000 million